Operations Report – March 27th

Crew 191 Operations Report 27/03/2018

SOL:3

Name of person filing report: Wataru Okamoto

Non-nominal systems: N/A

Notes on non-nominal systems: N/A

Generator (hours run):13hours

Solar SOC 96% (Before generator is run at night)

Diesel 25%

Propane 70psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) ? 7gallons.

Water (trailer) 60%

Water (static) 80%

Trailer to Static Pump used ? no

Water (loft) Static to Loft Pump used yes

Water Meter: 7:50 6 132778.1->pump 14 16:00 10 132790.2

Toilet tank emptied: yes

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) –

Oil Added –

ATV Fuel Used: – 350.2 45%->75%

Hours the ATVs were Used today: –

Notes on ATVs:-

Deimos rover used: –
Hours:
Beginning charge:
Ending charge:
Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours:
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: –
Hours:27.9h
Beginning charge:100%
Ending charge: 100%
Currently charging: 100%

Opportunity rover used: –
Hours: 26.9h
Beginning charge: 100%
Ending charge: 100%
Currently charging: 100%

Curiosity rover used: –
Hours:22.1h
Beginning charge: 100%
Ending charge: 100%
Currently charging: 100%

HabCar used and why, where? –

General notes and comments: -I didn’t turn off the Diesel generator in this morning. I turned off the Diesel generator after 50 minuites. It’s my mistake.I’ll turn off when amount of battery is over 80%.

Summary of internet: Under control

Summary of suits and radios:No.5 Backpack Battery was 6.8V after charging. cSo another 13V battery was installed.2 Backpack belt have repaired

Summary of Hab operations:No problem

Summary of GreenHab operations: No problem

Summary of ScienceDome operations: Only use checking the solar system.

Summary of RAMM operations: N/A

Summary of health and safety issues: No problem

Questions, concerns and requests to Mission Support:Please charge the Diesel gas.

Greenhab Repot – March 27th

GreenHab Report
Kai Takeda
27 MAR 2018

Environmental control:

Ambient with window/door open
Shade cloth on
Fan on
Heater on(morning)

Average temperatures: 22.8℃
Low 18.5℃
High 31.3℃

Hours of supplemental light: None

Changes to crops: Added hydroponic crops.

Daily water usage for crops: 2gal

Time(s) of watering for crops: 08:30 and 15:30

Moringa research observations: None

Changes to research plants: No

Aquaponics: None

Narrative: I set a Japanese hydroponic machine “foop” and my hydroponic crops.
I harvested tomatoes, green peppers, leaf lettuce and herb.

Support/supplies needed: None

EVA Report – March 27th

EVA Report

EVA #3 – 27 Mar 2018 – Sol3
Crew members: Fumiei Morisawa (EVA leader), Venzha Christ, Makoto Kawamura

Objectives:
– Rubbish collections around MDRS
– Lecture about GPS device handling
– Scouting for Venzha’s CMEs project

Operations:
9:00 – Entering the airlock
9:03 – Collecting an engineering log in EVA
9:34 – Start to garbage collecting and scouting
10:50 – Turn back to the HAB
11:26 – Enter the airlock
11:29 – EVA end

Narrative:
It is sunny and wind is calm today, a condition is nice for collecting rubbish. We have patrolled the foot of North Ridge and collected various rubbish. A piece of cloth, plastic beverage bottle, old glove, and so on. Some of the large object that we couldn’t carry by hand are leaved there as it was. So we need to recollect them on another EVA with a vehicles later on. All the rubbish have recorded by GPS coordinates and photos data.

Vehicles
– none

Mission Summary – Crew 190

 

It’s been twelve SOLs that Crew 190 landed on Mars. They rapidly settled down in the Hab and got to work. Quickly they learned to live with each other. A bit formally, we distinguish in this report five main activities that had been shared and managed during the time living in the station (see Figure 1): scientific work, team management, housework, EVAs and social activities.

SCIENTIFIC WORK

Scientific challenge was the main motivation of our stay at the MDRS. On a typical daily schedule, it usually primes on the rest of the activities. In this section, we provide a summary of the objectives and outcomes amongst the several experiments that were initially planned by the crew. Here we briefly present our different experiments.

Psychological and emotional aspects are paramount in this kind of mission. Martin Roumain, the Health and Safety Officer and biomedical researcher, evaluated the impact of confinement by monitoring short-term memory and reflexes throughout the mission. He also studied the accelerated degradation of drugs by the Martian environment using a spectrophotometer (Figure 2). This device has also been used by Maximilien Richald, chemist and Crew Commander of this mission. Maximilien focused on the chemical profile of Martian soil in view of an eventual use in agriculture. Food self-sufficiency being essential for long duration space missions, Mario Sundic, botanist and GreenHab Officer, has designed a vertical hydroponic system, reducing the water needs thanks to this closed circuit. Frédéric Peyrusson, Crew Biologist, tested the benefits of hydrogels on plants growth. Moreover, he studied the ability of known bacteria to resist to a harsh environment and figured out the biocompatibility of terrestrial life on Mars. Our second Biologist, Ariane Sablon (Figure 2), isolated fermentating bacteria from human saliva in view of making possible the preparation of sourdough bread in situ. Bastien Baix, the Crew Engineer, created a self-made 3D-map of the station and its surroundings using an aerial drone. Our physicist and Crew Astronomer Sophie Wuyckens, also contributed to terrain analysis by setting up a method based on cosmic radiations measurements (Figure 3). She was also in charge of the Musk Observatory. All these experiments needed to be perfectly coordinated. That was the role of Michael Saint-Guillain, the computer scientist and Executive Officer, who designed an algorithm that helped the scheduling of the various experiments conducted by the crew members.

Figure 3: Sophie Wuyckens working on her muon detector, inside the RAM module.

Besides the experiments, extravehicular activities (EVAs) required a significant amount of time. Of critical importance for some experiments, the EVAs also revealed a positive impact on the mental of the entire team and unique opportunities to contemplate the scenic martian landscape.

TEAM MANAGEMENT

Even though we stayed only a couple of weeks, such a mission requires sound organisation and systematical rescheduling. At the center of our eight people crew, the commandant Maximilien Richald had the hard responsibility of managing the entire team in a holistic way, dealing with all dimensions of the mission : experiments, housework, social behavior… Which had to be discussed during daily team meeting (Figure 4).

Coordination of the scientific operations (including EVAs, manipulations in the ScienceDome, solar observations, homework) had been closely monitored by Michael Saint-Guillain. At the end of each day, just before the CapCom, the scientific outcomes were used as input to the scheduling algorithm which was then used to recompute a schedule for the rest of the mission.

HOUSEWORKING

As part of a large team enclosed in a quite small living space, we all recognized the critical importance of well-balanced houseworking. The time required for meal preparation should not be underestimated, as we are cooking with unusual freeze-dried ingredients, which, by the way, did not prevent us from cooking some masterpieces (Figure 5) !

 

SOCIAL ACTIVITIES

Last but not least, it is worth to mention that social activities definititly contributed to the success of the mission. Even with the best team composition, maintaining a good mood is not trivial and has a significant impact on the global outcomes of the mission. Fortunately, the Mars Society staffed the MDRS with a few interesting games, as the one we played on Figure 6 (left). On the right in Figure 6, we even observe a singular birthday event, quite uncommon on Mars !

Sol Summary – March 25th

Hi CapCom,
here follows our 1st Sol Summary Report:

Crew 191 Sol Summary Report
25 Mar 2018

Sol 01

Summary Title:
Preparing for Simulation and live broadcasting in Japan
Miho Tsukishiro (XO)

Mission Status:
Preparing our simulation from sol 02

Sol Activity Summary:
8:00 Breakfast
8:40 Briefing
9:30 Spacesuits & how to go outside during EVA lecture by commander
10:40 preparing for simulation / each
12:10 Lunch – crew meeting
12:40 preparing for simulation / each
16:30 live broadcasting by Japanese TV
17:00 lecture about environment of MDRS with ATV (commander & engineer)
18:00 gathering Reports
19:00 Capcom

Look Ahead Plan:
Our simulation starts by 12:00AM 26 March 2018
The first EVA plan: our commander gives us lecture about ATVs

Anomalies in work:
N/A

Weather:
Cloudy; Very strong winds day and night.

Crew Physical Status:
Good

EVA:
N/A

Reports to be file:
– Commander and Journalist Greetings
– SOL Summary Report
– SOL Essay
– Engineer Report
– GreenHab Report
– 8 Pictures
– EVA Request

Support Requested:
N/A

Best,
Miho

EVA Report – March 22nd

EVA Report

EVA #12 – 22th March 2018 – Sol 11

Crew members: Bastien Baix (EVA leader), Michael Saint-Guillain, Mario Sundic, Sophie Wuyckens

Objectives:
– Drone mapping and drone filming by Bastien
– Photos by Mario
– Sophie will put her detector outside of the RAM

Operations:
9:43 – Entering the airlock
9:47 – Leaving the MDRS with 2 rovers
9:55 – Arriving at Zubrin’s Head
10:20 – We start climbing a high hill
10:42 – Top of the mountain
10:50 – Bastien map the mountain
11:05 – We go down of the mountain
11:40 – We leave Zubrin’s Head
11:50 – Arrival at the MDRS, Mario and Bastien go inside, Michael and Sophie install the muon detector 12:05 – Michael and Sophie enter the airlock

Narrative:
One of our last EVA took place on soil 12 at 9:40. I am Bastien Baix and I was EVA leader today. We decided to climb a high hill today. To do so, we went to Zubrin’s Head. Near this location, we located our objective and we start our climbing. It was particularly difficult as the slope was nearly vertical! As Sophie is relatively clumsy, we made her a safety line…

This climbing allowed us to take amazing pics and vids with the drone. Once we made it to the top, the view was amazing. We’ve seen miles around.

When we decided to go back and reach our rovers, we couldn’t find them anymore! We pass half an hour to finally find our way back to them and ride to the MDRS. It was an awesome but sweaty EVA!

Vehicles:
Curiosity and Opportunity rovers

Operations Report – March 22nd

SOL: 11

Name of person filing report: Bastien BAIX
Non-nominal systems: noisy pump, see end of report below
Notes on non-nominal systems: –

Generator (hours run): Turned off at 12.40 pm and turned on at 6.30 pm Solar – SOC 80% (Before generator is run at night)
Diesel – ~25%
Propane – 80%
Ethanol Free Gasoline (5 Gallon containers for ATV) – 1 ½ full jerrycans.

Water (trailer) – 0 gallons.
Water (static) – ~500 gallons
Water (third tank) – 0 gallons
Trailer to Static Pump used – NO
Water (loft) – Static to Loft Pump used – YES
Water Meter: 132 585.7
Toilet tank emptied: NO

ATV’s Used: –
Oil Added: NO
ATV Fuel Used: NO
Hours the ATVs were used today: _
Notes on ATVs: –

Deimos rover used: NO
Hours: 116.7
Beginning charge: 100%
Ending charge: –
Currently charging: YES

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours:
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: NO
Hours: 26.9
Beginning charge: 100%
Ending charge: –
Currently charging: NO

Opportunity rover used: YES
Hours: 26
Beginning charge: 100%
Ending charge: 93%
Currently charging: NO

Curiosity rover used: YES
Hours: 20.8
Beginning charge: 100%
Ending charge: 91%
Currently charging: YES

HabCar used and why, where: –
General notes and comments: –
Summary of internet: ~ 120 Mb remaining
Summary of suits and radios: –
Summary of Hab operations: –
Summary of GreenHab operations: –
Summary of ScienceDome operations: –
Summary of RAMM operations: –
Questions, concerns and requests to Mission Support:

Update on the pump: I made pictures of the pump (and the old pump too). – Model: Pentair Shurflo
– 115 VAC
– 60Hz
– 3,3 GPM /12,5 LPM of max flow
– 104 Watts
– Pressure: 45 PSI / 3,1 bar

Sol Summary – March 22nd

Crew 190 Sol Summary Report
22 March 2018

Sol 11

Summary Title:
Happy Birthday!

Michael Saint-Guillain (XO)

Mission Status:
Ready to continue simulation for the last day on Mars (Sol 12), 08:00

Sol Activity Summary:
8:20 Gym
8:35 Medic inspection
8:50 Breakfast
9:40 EVA: Bastien (cartography), Sophie (placement of muon detector), Mario (photos), Michael (support) 12:05 End of EVA – Crew meeting – Relaxation
13:00 Lunch
14:00 Scientific & academic work – Interviews continued (Ariane) 18:00 Report redaction
19:00 CapCom

Look Ahead Plan:
Sol 12:
Notes: Last EVA for Maximilien & Ariane (soil sampling), Fred (collection of the experiment he placed outside several days ago), Martin (support). Part of the crew will conclude their experiments in the afternoon while the others will do some clean up inside the MDRS. Sol 12 will be the last day under simulation. On Sol 13 we will welcome Crew 191.

Anomalies in work:
None.

Weather:
Very cloudy all day long… Preventing from sun telescope activities, despite the sun storm announced. Too bad. Hopefully tomorrow will still allow us to observe residual sun activities.

Crew Physical Status:
Still tired, but we observe an improvement compared to the couple of last days! However we have significantly more difficulty to recover after the morning gym.

EVA:
See EVA report.

Reports to be filed:
* Commander report
* Journalist report
* EVA#12 report
* EVA#13 request for Sol12
* Engineer report
* Pictures
* Summary report

Support Requested:
None

Commander Report – March 22nd

Crew 190 – sol11 – crew commander report 22-03-2018

First, we made a short gym session for waking up. Then, we ate breakfast and we wished a happy birthday to Michael who is now 31 years old. Sophie, Bastien, Mario and Michael left for an EVA up to “Zubrin Head” place. They took some pictures and Bastien mapped the area. Afterwards, they installed the muons detector outside of the station before coming back inside the Hab.

At midday, Ariane started to cook the lunch and we spent time eating and speaking for Michael Birthday’s. To complete our afternoon, everyone worked to his scientific project. Last manipulations will be done tomorrow morning.
Tonight we will have dinner followed by a cake kindly prepared by the crew commander especially for Mick’s birthday.

Tomorrow, we plan to end the experiments and have a small EVA before starting to clean up the whole station. Our way back to Earth is now in the mind of each crew members.

Regarding the station, nothing to declare.